Find four consecutive odd numbers such that the sum of the first and last is 64.

thank you

Found 2 solutions by stanbon, RAY100:
Answer by stanbon(75887)

Find four consecutive odd numbers such that the sum of the first and last is 64.
------------------
1st: 2x-3
2nd: 2x-1
3rd: 2x+1
4th: 2x+3
--------------------
(2x-3)+(2x+3) = 64
4x = 64
x = 16
-------------
1st: 2x-3 = 29
2nd: 2x-1 = 31
3rd: 2x+1 = 33
4th: 2x+3 = 35
=======================
Cheers,
Stan H.



Answer by RAY100(1637)

Remember that odd numbers are 2 digits apart
.
let x =1 st number
.
let x+2 = 2nd number
.
let x+4 = 3 rd number
.
let x+6 = 4th number
.
Sum of 1st and 4th is,,,(x) + (x+6) = 64
.
2x +6 =64
2x = 58
x=29, 1st number
4 th number is x+6 = 29 +6 = 35
.
checking, 29 +35 =64,,,and 29,31,33,35 is valid